OpenStudy (anonymous):

Suppose that a large pump can empty a swimming pool in 50 hours and a small pump can empty the pool in 80 hours. How long will it take to empty the pool if both pumps are used?

OpenStudy (anonymous):

\[\frac{1}{50}+\frac{1}{80}=\frac{1}{x} \]Solve the above for x, the total time for both pumps working together.\[\frac{400}{13}=30.7692 \]